OpenStudy (anonymous):

Solve the equation and write complex solutions in standard form. 3t^2+45=-3

OpenStudy (solomonzelman):

3t^2+45=-3 3t^2+45-45=-3-45 3t^2=-48 t^2=-16 t= +- 4i

OpenStudy (solomonzelman):

t= 4i, -4i

OpenStudy (anonymous):

Great! Thank you!!!

OpenStudy (solomonzelman):

Anytime!